OpenStudy (anonymous):

Subtract and simplify: (2x^2 + 5x – 4) – (3x^2 – 7x + 1)

OpenStudy (anonymous):

answers: A. –x^2 + 12x – 5 B. –x^2 – 2x – 3 C. 5x^2 – 2x – 3 D. x^2 + 12x – 3

OpenStudy (bahrom7893):

2x^2-3x^2 + 5x+7x - 4-1

OpenStudy (anonymous):

-x^2+12x-3

OpenStudy (anonymous):

oh so its D

OpenStudy (bahrom7893):

read what you have for D and then try again.

OpenStudy (anonymous):

oh so its A

OpenStudy (bahrom7893):

@Jhannybean i think she needs your help. @desiraedallas1999 Try again.

OpenStudy (anonymous):

x^2+12x-3?

OpenStudy (anonymous):

idk i thought was doing it right 2x^2-3x^2 + 5x+7x - 4-1 1x^2+12x-3

OpenStudy (bahrom7893):

what's -4-1?

OpenStudy (anonymous):

-4-1= -4+-1= -3

OpenStudy (bahrom7893):

-4-1 = -(4+1) = -5

OpenStudy (anonymous):

oh so A right: -x^2+12x-5

OpenStudy (bahrom7893):

yea

OpenStudy (anonymous):

oh ok thanks :)